# https://aimotive.com/

### v2

@@ Abstract
Keywords:
self-driving – automatic annotation – neural networks – 3D localization
@@ 1 Introduction
Deep learning models currently used in self-driving cars require a vast amount of training data to ensure accurate predictions in all scenarios. As a consequence, there is a need for labeling every dynamic and static object with 3D bounding boxes and additional attributes over hundreds or thousands of hours of driving. However, manually creating these labels is expensive, time-consuming, and error-prone. While several datasets with 3D bounding box annotations are available for dynamic objects [1], [2], [3], [4], the number of available static object datasets with 3D annotations [5], especially those containing distant objects, is remarkably limited. As a result, there is a significant interest in automating the generation of such training data without human intervention. Our primary goal is to provide accurate 3D bounding boxes for traffic management objects, ensuring that the projected 2D bounding boxes in the camera image encompass objects from a wide range of viewing angles and distances. This step is crucial for all downstream tasks of the proposed method, such as classification or optical character recognition. Since the data recording process typically involves multiple sensors and a high frame rate, this requirement is easily met.
The main contribution of this work is a novel method that provides accurate positioning with an average mean distance of 0.2-0.3 meters and temporally consistent 3D bounding boxes of traffic management objects from up to 200 meters away. Our method also determines additional attributes such as traffic light state, traffic light mask type, traffic sign type, and occlusion. The proposed solution is simple yet effective, relying solely on 2D images and Global Navigation Satellite System/Inertial Navigation System (GNSS/INS) data, without the need for expensive active sensors like LiDAR. Furthermore, we publish a representative dataset [^1] https://github.com/aimotive/aimotive_tl_ts_dataset, automatically generated using our algorithm, under a CC BY-NC-SA 4.0 license, allowing the research community to use it for non-commercial research purposes. To our knowledge, no publicly available large-scale dataset including distant objects currently exists that contains accurate 3D bounding boxes of traffic management objects, particularly traffic lights.
@@ 5.1 Validation Challenges
Precise localization of traffic management objects on a large scale is extremely challenging due to issues such as sensor limitations described in Section 2. This challenge explains why there is still no publicly available dataset with long-range 3D annotations for traffic signs and traffic lights. Although Mapillary provides global latitude and longitude coordinates for traffic signs, the accuracy is low, and there is no information about the vertical position, extent, or orientation to accurately place these objects in the local coordinate system of a driving scene. Popular autonomous driving datasets like nuScenes, KITTI, and Waymo present additional challenges. Among these, only Waymo provides 3D bounding boxes for traffic signs, but it lacks GNSS information for the camera frames, which is necessary to evaluate our algorithm on a dataset. Moreover, we are not aware of any publicly available traffic light datasets with 3D annotations, especially those containing distant objects. Given these difficulties, we have decided to validate our algorithm using manually annotated in-house benchmark datasets.
5.2 Automatic Traffic Sign Annotation
@@
We validated the traffic sign automatic annotation performance on a 7-kilometer route in San José, California, which included both highway and urban sections (see the validation route in Figure 8). In total, 183 traffic signs were manually annotated with oriented 3D bounding boxes using LiDAR point cloud data. This manually created map was projected into the instantaneous coordinate systems of the vehicle, allowing for a detailed comparison with the automatic annotation. All metrics were calculated within the range of [-10m, 10m] lateral and [0m, 200m] longitudinal positions of the instantaneous coordinate system. The association distance threshold was set to 1 meter, and we calculated localization precision and recall related to the bounding box center. The automatic annotation method achieved 97.08% precision and 95.33% recall (see Table 1 for more detailed results). It is worth noting that the lower recall value resulted from only six missed traffic signs on the highway section, which included traffic signs with categories less relevant for self-driving (e.g. destination distance, interchange advance exit).
We also evaluated the localization errors of true positive detections using the absolute mean distance between the 3D bounding box centers and the annotations. Moreover, the absolute orientation error of the annotations is also evaluated. Our algorithm achieves low localization (0.3 meters) and orientation (11.09 degrees) errors.
@@
Table 1: Quantitative evaluation results of our automatic annotation method for traffic signs.
@@
5.3 Automatic Traffic Light Annotation
We validated the automatic traffic light annotation algorithm at several intersections in Palo Alto, California. The validation route is approximately 1.3 kilometers long and includes 40 traffic lights (see the validation route in Figure 9). The 3D bounding boxes of the traffic lights, as well as their states, were manually annotated. Consequently, we measured both localization performance and traffic light state classification accuracy. In the association metrics, a true positive means the prediction is within 1 meter of the ground truth and the predicted class is correct. All metrics were calculated within the range of [-10m, 10m] lateral and [0m, 200m] longitudinal positions of the instantaneous coordinate system. Our method achieved 91.13% precision and 95.87% recall. The absolute localization error between the bounding box centers is 22 centimeters, and the orientation absolute error is 10.49 degrees. The traffic light color state classification accuracy is 94%.
@@ 5.3 Automatic Traffic Light Annotation
Table 2: Quantitative evaluation results of our automatic annotation method for traffic lights.
@@ 6 Conclusion
Despite self-driving developments that have been conducted for several decades, there is still no publicly available large-scale dataset with 3D annotated traffic lights and traffic signs. This indicates that annotating traffic management objects is challenging, even with manual resources. This is especially true for traffic lights, which are difficult to detect in LiDAR point clouds even for humans, as their physical characteristics (e.g., small size, high placement, and black coating) make it challenging for the sensor to produce easily detectable reflections. In this work, we developed a fully automated method to generate temporally consistent 3D bounding boxes with high localization precision for traffic lights and traffic signs, which can be used to train image-based perception models for self-driving cars. Additionally, we released a public dataset generated by our algorithm, available under a CC BY-NC-SA 4.0 license, allowing the research community to use it for non-commercial research purposes.

### v4

# Accurate Automatic 3D Annotation of Traffic Lights and Signs for Autonomous Driving

@@ 1 Introduction
Deep learning models currently used in self-driving cars require a vast amount of training data to ensure accurate predictions in all scenarios. As a consequence, there is a need to label every dynamic and static object with 3D bounding boxes and additional attributes over hundreds or thousands of hours of driving. However, manually creating these labels is expensive, time-consuming, and error-prone. While several datasets with 3D bounding box annotations are available for dynamic objects [14], [1], [7], [9], the number of available static object datasets with 3D annotations [5], especially those containing distant objects, is remarkably limited. As a result, there is a significant interest in automating the generation of such training data without human intervention. Our primary goal is to provide accurate 3D bounding boxes for traffic management objects, ensuring that the projected 2D bounding boxes in the camera image encompass objects from a wide range of viewing angles and distances. This step is crucial for all downstream tasks of the proposed method, such as classification or optical character recognition. Since the data recording process typically involves multiple sensors and a high frame rate, this requirement is easily met.
The main contribution of this work is a novel method that provides accurate positioning with an average mean distance of 0.2-0.3 meters and temporally consistent 3D bounding boxes of traffic management objects up to 200 meters away. Our method also determines additional attributes such as traffic light state, traffic light mask type, traffic sign type, and occlusion. The proposed solution is simple yet effective, relying solely on 2D images and Global Navigation Satellite System/Inertial Navigation System (GNSS/INS) data, without the need for expensive active sensors like LiDAR. Furthermore, we publish a representative dataset, automatically generated using our algorithm, under a CC BY-NC-SA 4.0 license, allowing the research community to use it for non-commercial research purposes [^1] https://github.com/aimotive/aimotive_tl_ts_dataset. To our knowledge, no publicly available large-scale dataset including distant objects currently exists that contains accurate 3D bounding boxes of traffic management objects, particularly traffic lights.
@@ 5.1 Validation Challenges
Precise localization of traffic management objects on a large scale is extremely challenging due to issues such as sensor limitations described in Section 2. This challenge explains why there is still no publicly available dataset with long-range 3D annotations for traffic signs and traffic lights. Although Mapillary provides global latitude and longitude coordinates for traffic signs, the accuracy is low, and there is no information about the vertical position, extent, or orientation to accurately place these objects in the local coordinate system of a driving scene. Popular autonomous driving datasets like nuScenes, KITTI, and Waymo present additional challenges. Among these, only Waymo [15] provides 3D bounding boxes for traffic signs and has GNSS information for the camera frames, which is necessary to evaluate our algorithm on a dataset. However, it contains annotations up to only 77-78 meters from the observer, and there is no information about the relevance of the traffic sign to the ego vehicle, hence we cannot directly measure precision or recall, but only a distance error between the associated ground truth-prediction pairs. Moreover, we are unaware of publicly available traffic light datasets with 3D annotations, especially those containing distant objects. Given these difficulties, we have decided to validate our algorithm not only on the Waymo traffic sign dataset but also using manually annotated in-house benchmark datasets.
Figure 8: Qualitative comparison of Waymo ground truth (green) and auto-annotated (red) 3D bounding boxes (Left: segment-14811410906788672189_373_113_393_113_with_camera_labels; Right: segment-10203656353524179475_7625_000_7645_000_with_camera_labels). The annotated traffic sign types in the ground truth and in the automatic annotation can be very different.
Figure 9: Evaluation results (in 4 m x 10 m blocks until 80 meters) of the proposed algorithm measured on all traffic sign boxes related to the Waymo validation set (Left: Mean error in bounding box center estimation (0.32 meters). Right: Mean absolute error in box orientation (12.31 degrees). (best viewed by zooming in)
5.2 Validation of the method on the Waymo dataset
[…]
Table 1: Quantitative evaluation results of our automatic annotation method for all traffic signs of the Waymo validation dataset.
Metric Result
[…]
Orientation error 12.31 degrees
Figure 10: Evaluation results (in 4 m x 10 m blocks until 80 meters) of the proposed algorithm measured on the manually selected speed limit and stop signs related to the Waymo validation set. Left: Mean error in bounding box center estimation (0.28 meters). Right: Mean absolute error in box orientation (8.78 degrees).
Table 2: Quantitative evaluation results of our automatic annotation method for speed limit and stop signs of the Waymo validation dataset.
Metric Result
Recall 93.76 %
[…]
Orientation error 8.78 degrees
[…]
5.3 Validation of Automatic Traffic Sign Annotation on in-house dataset
@@
We also validated the traffic sign automatic annotation performance on a 7-kilometer route in San José, California, which included both highway and urban sections (see the validation route in Figure 11). In total, 183 traffic signs were manually annotated with oriented 3D bounding boxes using LiDAR point cloud data. This manually created map was projected into the instantaneous coordinate systems of the vehicle, allowing for a detailed comparison with the automatic annotation. All metrics were calculated within the range of [-10m, 10m] lateral and [0m, 200m] longitudinal positions of the instantaneous coordinate system. The association distance threshold was set to 1 meter, and we calculated localization precision and recall related to the bounding box center. The automatic annotation method achieved 97.08% precision and 95.33% recall (see Table 3 for more detailed results). It is worth noting that the lower recall value resulted from only six missed traffic signs on the highway section, which included traffic signs with categories less relevant for self-driving (e.g. destination distance, interchange advance exit).
[…]
@@
Figure 12: Evaluation results (in 4 m x 10 m blocks until 200 meters) of the proposed algorithm measured on our manually annotated in-house traffic sign dataset. Left: Mean error in bounding box center estimation (0.3 meters). Right: Mean absolute error in box orientation (11.09 degrees).
Figure 13: Precision and recall (in 4 m x 10 m blocks until 200 meters) of the proposed algorithm measured on our manually annotated in-house traffic sign dataset. Left: Precision (97.08 %). Right: Recall (95.33 %).
Table 3: Quantitative evaluation results of our automatic annotation method for traffic signs on in-house dataset.
@@
5.4 Validation of Automatic Traffic Light Annotation on in-house dataset
We validated the automatic traffic light annotation algorithm at several intersections in Palo Alto, California. The validation route is approximately 1.3 kilometers long and includes 40 traffic lights (see the validation route in Figure 14). The 3D bounding boxes of the traffic lights, as well as their states, were manually annotated. Consequently, we measured both localization performance and traffic light state classification accuracy. In the association metrics, a true positive means the prediction is within 1 meter of the ground truth and the predicted class is correct. All metrics were calculated within the range of [-10m, 10m] lateral and [0m, 200m] longitudinal positions of the instantaneous coordinate system. Our method achieved 91.13% precision and 95.87% recall. The absolute localization error between the bounding box centers is 22 centimeters, and the orientation absolute error is 10.49 degrees. The traffic light color state classification accuracy is 94%. Detailed metrics are shown in Fig. 15 and Fig. 16.
@@ 5.4 Validation of Automatic Traffic Light Annotation on in-house dataset
Figure 15: Evaluation results (in 4 m x 10 m blocks until 200 meters) of the proposed algorithm measured on our manually annotated in-house traffic light dataset. Left: Mean error in bounding box center estimation (0.22 meters). Right: Mean absolute error in box orientation (10.49 degrees).
Figure 16: Precision and recall (in 4 m x 10 m blocks until 200 meters) of the proposed algorithm measured on our manually annotated in-house traffic light dataset. Left: Precision (91.13 %). Right: Recall (95.87 %).
Table 4: Quantitative evaluation results of our automatic annotation method for traffic lights on in-house dataset.
@@ 6 Conclusion
Despite self-driving developments that have been conducted for several decades, there is still no publicly available large-scale dataset with 3D annotated traffic lights and traffic signs. This indicates that annotating traffic management objects is challenging, even with manual resources. This is especially true for traffic lights, which are difficult to detect in LiDAR point clouds even for humans, as their physical characteristics (e.g., small size, high placement, and black coating) make it challenging for the sensor to produce easily detectable reflections. In this work, we developed a fully automated method to generate temporally consistent 3D bounding boxes with high localization precision for traffic lights and traffic signs, which can be used to train image-based perception models for self-driving cars. Additionally, we released a public dataset generated by our algorithm, available under a CC BY-NC-SA 4.0 license, allowing the research community to use it for non-commercial research purposes [^3] https://github.com/aimotive/aimotive_tl_ts_dataset.